# How Combined Pairwise and Higher-Order Interactions Shape Transient Dynamics

### v1

@@ Abstract
Understanding how species interactions shape biodiversity is a core challenge in ecology. While much focus has been on long-term stability, there is rising interest in transient dynamics—the short-lived periods when ecosystems respond to disturbances and adjust toward stability. These transitions are crucial for predicting ecosystem reactions and guiding effective conservation. Our study introduces a model that blends pairwise and higher-order interactions, offering a more realistic view of natural ecosystems. We find pairwise interactions slow the journey to stability, while higher-order interactions speed it up. This model provides fresh insights into ecosystem resilience and recovery, helping improve strategies for managing species and ecological disruptions.
@@ I Introduction
However, studying these short-term dynamics presents challenges because current mathematical theories are not yet well-developed, making predictions and management more difficult. Nevertheless, research is increasingly focused on understanding the causes and impacts of transient dynamics and finding ways to manage them. Examples include disease outbreaks Tao et al. (2021), shifts in species populations, and significant changes in ecosystems, such as coral reef degradation Norström et al. (2009). Early studies have also explored related concepts, such as cyclical plant succession and species’ behavior in lakes and savannas under varying conditions Van Langevelde et al. (2003); Scheffer et al. (1993). Recognizing the importance of transient dynamics is crucial because these temporary periods can lead to unexpected changes in ecosystems that might not be predicted by focusing solely on long-term behavior Hastings et al. (2018); Morozov et al. (2020). Understanding these dynamics is especially important when managing invasive species and in conservation and restoration efforts, where setting realistic recovery timelines and identifying critical points for intervention are essential. Additionally, transient dynamics reveal complex interactions between species, such as predator-prey relationships and competition, that may be less apparent once the ecosystem reaches a stable state. The importance and complexity of transient dynamics in physical systems have also recently gained attention. There have been reports of transient chaotic behavior in dynamical systems Dudkowski et al. (2022), even in cases where the attractor is hidden within the phase space Nag Chowdhury and Ghosh (2020). In fact, it may be possible that a dynamical system may become temporarily trapped at a local maximum or minimum within a potential landscape for a long time, giving the impression of stability. However, a minor disturbance—such as noise or an external force—can enable the system to overcome this barrier and transition to the lowest energy (or most stable) configuration Koch et al. (2024); Chowdhury et al. (2020); Bovier and Den Hollander (2016); Kelso (2012).
On the other hand, traditionally, models have focused on either pairwise interactions Nag Chowdhury et al. (2024, 2023a, 2019a) or higher-order interactions (HOI) Battiston et al. (2020); Gibbs et al. (2022); Zhang et al. (2023); Battiston et al. (2021) exclusively. However, real-world ecosystems are likely to exhibit a mix of these interaction types, with species interactions containing varying fractions of higher-order components. Such mixed-interaction frameworks are essential to better reflect the complexities of natural systems and provide more accurate insights into their dynamics. Recent research has focused on identifying HOIs from ecological data and making HOIs more common in ecological systems Terry et al. (2020); Li et al. (2021); Raj et al. (2024); Kleinhesselink et al. (2022); Barbosa et al. (2023).
@@
In classical niche theory Chase and Leibold (2009); Soberón and Nakamura (2009), species coexistence within the same niche is often challenged by competitive exclusion, leading weaker species to be out-competed by stronger ones. Conversely, neutral theory Brokaw and Busing (2000) posits that species are ecologically equivalent, with biodiversity emerging from a stochastic balance between speciation and extinction. While both theories have significantly advanced our understanding, they also have notable limitations. For instance, niche theory struggles with the paradox of the plankton, where the number of coexisting species exceeds the number of available limiting resources Huisman and Weissing (1999). Similarly, neutral theory’s “neutral drift” is incompatible with the observed stability in forest diversity Clark and McLachlan (2003).
Recent developments in ecological modeling have introduced game theory-based approaches Nag Chowdhury et al. (2023b, 2021a); Roy et al. (2022, 2023); Ahmadi et al. (2023); Nag Chowdhury et al. (2021b), such as intransitive cyclic competition models like the rock-paper-scissors (RPS) game, which maintain biodiversity through cyclical dominance. These models demonstrate that even in hierarchical systems, intransitive competition can stabilize species coexistence, a phenomenon known as the “stabilizing effect of intransitivities.” Empirical evidence supports the prevalence of intransitive competition in various ecological communities, such as bacterial strains Kirkup and Riley (2004); Laird and Schamp (2006); Hibbing et al. (2010); Szolnoki et al. (2014); Park et al. (2017); Bhattacharyya et al. (2020); Islam et al. (2022); Chatterjee et al. (2023) and phytoplankton Soliveres et al. (2015) and parasite-grass-forb Cameron et al. (2009).
@@
Although research on systems that consider both pairwise and higher-order interactions is still emerging, it is gaining attention Gibbs et al. (2022); Malizia et al. (2024); Kundu and Ghosh (2022). One recent study Van Giel et al. (2024) examined how the speed at which higher-order interactions emerge affects the stability and evolution of ecological networks, offering new insights into this underexplored area. Another study Gibbs et al. (2024) found that while higher-order interactions can lead to equilibrium in species abundance, they do not always ensure stable coexistence. However, when weak or cooperative pairwise interactions align with higher-order interactions, they can foster robust coexistence in diverse ecosystems. Using the generalized Lotka-Volterra model, Ref. Singh and Baruah (2021) derive a rule showing that while negative higher-order interactions can stabilize species coexistence by strengthening intraspecific competition, positive higher-order interactions can do so across a wider range of conditions by alleviating pairwise competition, and their results extend to multispecies communities, emphasizing the role of negative intraspecific HOIs in maintaining diversity. Our focus differs as we seek to better understand transient dynamics and explore how they respond to the simultaneous presence of both pairwise and higher-order interactions.
@@
Figure 1: Dynamics of model systems for (a) Pairwise $(\gamma_{1}=1.0)$ (b) Non-pairwise $(\gamma_{2}=1.0)$ (c) Mixed system $(\gamma_{1}\neq 0$, $\gamma_{2}\neq 0$, $\gamma_{1}+\gamma_{2}=1.0)$. The initial condition $(0.45,0.3,0.25)$ is chosen randomly within the $(0,1)\times(0,1)\times(0,1)$ maintaining the constraint $x_{1}(0)+x_{2}(0)+x_{3}(0)=1.0$. As shown in subfigure (a), the pairwise interaction system does not converge to a steady state. However, introducing higher-order interactions significantly alters this behavior. Interestingly, the system with both pairwise and higher-order interactions converges more slowly compared to the system with only higher-order interactions, although both ultimately reach the same equilibrium point $\big{(}\frac{1}{3},\frac{1}{3},\frac{1}{3}\big{)}$.
@@ II.1 Description
where $B_{ij}^{1}=A_{ij}-A_{ji}$, and $B_{ijk}^{2}=2A_{ij}A_{ik}-A_{ji}A_{jk}-A_{ki}A_{kj}\;\forall i,j=1,2,3,\cdots,N$.
Here, $\gamma_{i}\in(0,1)$ represents the contribution factor, which indicates the percentage of $B^{i}$ that contributes to the total interaction. So,
@@
$A$ matrix encodes information about interactions between two species. As there is $N$ species, $A$ is an $N\times N$ matrix. $A_{jk}$ term account for the winning probability of $j$-th species over the $k$-th species. Hence,
@@ II.2 Solution and Global Stability Analysis
Figure 2: Dynamics of mixed systems with $N=3$ for different values of $\gamma_{1}$ and $\gamma_{2}$: (a) $\gamma_{1}=0.7,\gamma_{2}=0.3$, (b) $\gamma_{1}=0.5,\gamma_{2}=0.5$, (c) $\gamma_{1}=0.3,\gamma_{2}=0.7$. Starting from the initial condition of $(0.45,0.3,0.25)$, as in Fig. 1, the system consistently converges to the coexistence equilibrium point, regardless of the relative contributions of pairwise and higher-order interactions. However, as the proportion of pairwise interactions decreases (i.e., as $\gamma_{1}$ diminishes), the transient time significantly shortens. This demonstrates that higher-order interactions play a crucial role in accelerating convergence towards equilibrium.
@@ II.2.1 Pairwise System
Now, we take the following Lyapunov function,
@@
Since $\sum_{j=1}^{N}A_{ij}x_{j}^{*}=\frac{1}{2}$, we have
@@ II.2.3 Mixed System
$-\gamma_{2}\big{[}2\sum_{i=1}^{N}x_{i}^{*}\sum_{j=1}^{N}\sum_{k=1}^{N}\big{(}A_{ij}A_{ik}+A_{ij}A_{jk}\big{)}x_{j}x_{k}+1\big{]}$
Now, using the same simplification as mentioned in the above two analyses, we finally get,
$\frac{dV}{dt}=-\gamma_{1}\big{[}0\big{]}-\gamma_{2}\big{[}-2\sum_{i=1}^{N}x_{i}^{*}\bigg{(}\sum_{j=1}^{N}A_{ij}\eta_{j}\bigg{)}^{2}\big{]}<0.$
@@ II.3 Transient dynamics
Figure 3: The boxplot illustrates how transient times vary with different values of $\gamma_{2}$ across various initial conditions. Mean transient time is indicated by orange dots and median is denoted by yellow line. For the considered interaction matrix $A$, we use ecologically relevant initial conditions and plot the transient times across varying levels of higher-order interactions. As $\gamma_{2}$ increases from $0.05$ (indicating a minimal presence of higher-order interactions) to $1.0$ (representing a system solely dominated by higher-order interactions), the transient times steadily decrease. While the higher values of $\gamma_{2}$ may appear to have minimal variation in transient times due to the large range of the y-axis, the inset provides a clearer view, showing that transient times with different initial conditions are actually well-distributed across this range.
Figure 4: (a) Mean transient time plotted against $\gamma_{2}$, with an exponential decay fit (red dashed line) illustrating how transient time decreases as $\gamma_{2}$ increases. (b) The most negative real part of the eigenvalues plotted against $\gamma_{2}$ (blue dots), with a linear fit (red dashed line). The negative real part of the leading eigenvalue in (b) provides a direct measure of convergence speed. The linear correlation between the most negative eigenvalue and transient time illustrates that greater negativity in the leading eigenvalue is associated with faster convergence to the same coexistence equilibrium point. This highlights the role of higher-order interactions in accelerating system stabilization and reducing transient oscillations.
@@ II.3.2 Role of Initial Abundances in Systems with a Fixed Interaction Matrix
In Figure 3, transient times are summarized using a boxplot, with the median indicated by the yellow line and the mean by the orange dot. Connecting lines between adjacent mean values provide a clearer view of the trend. As $\gamma_{2}$ increases, indicating a higher proportion of higher-order interactions, transient times decrease and variance reduces. This suggests that ecosystems with more complex interactions can stabilize more rapidly, which is crucial for predicting how real-world ecosystems maintain biodiversity. An inset in the figure highlights the reduction in transient time, showing how quickly systems approach stability with higher-order interactions. To create this figure, we selected $20$ different random initial conditions, each constrained by $x_{1}(0)+x_{2}(0)+x_{3}(0)=1.0$. We measured the transient time for each condition—defined as the period until density oscillations fall below $10^{-6}$ and then calculated the average of these transient times to obtain the mean value.
When $\gamma_{2}=0$, the system’s transient time is effectively infinite, leading to a rapid decline near zero and a more gradual decrease as $\gamma_{2}$ increases. To quantify this relationship, we fitted the mean transient time for $\gamma_{2}$ values ranging from $0.2$ to $1.0$ with an exponential distribution in Fig. 4(a). The resulting rate parameter of $2.27$ provides a statistical measure of how swiftly the system’s transient time diminishes as the proportion of higher-order interactions increases. This finding highlights the significant impact of higher-order interactions on accelerating the system’s convergence to equilibrium.

### v2

@@ Abstract
Understanding how species interactions shape biodiversity is a core challenge in ecology. While much focus has been on long-term stability, there is rising interest in transient dynamics—the short-lived periods when ecosystems respond to disturbances and adjust toward stability. These transitions are crucial for predicting ecosystem reactions and guiding effective conservation. Our study introduces a model that uses convex combinations to blend pairwise and higher-order interactions, offering a more realistic view of natural ecosystems. We find pairwise interactions slow the journey to stability, while higher-order interactions speed it up. Employing global stability analysis and numerical simulations, we establish that as the proportion of higher-order interactions (HOIs) increases, mean transient times exhibit a significant reduction, thereby underscoring the essential role of HOIs in enhancing biodiversity stabilization. Our results reveal a robust correlation between the most negative real part of the eigenvalues of the Jacobian matrix associated with the linearized system at the coexistence equilibrium and the mean transient times. This indicates that a more negative leading eigenvalue correlates with accelerated convergence to stable coexistence abundances. This insight is vital for comprehending ecosystem resilience and recovery, emphasizing the key role of HOIs in promoting stabilization. Amid growing interest in transient dynamics and its implications for biodiversity and ecological stability, our study enhances the understanding of how species interactions affect both transient and long-term ecosystem behavior. By addressing a critical gap in ecological theory and offering a practical framework for ecosystem management, our work advances knowledge of transient dynamics, ultimately informing effective conservation strategies.
@@ I Introduction
However, studying these short-term dynamics presents challenges because current mathematical theories are not yet well-developed, making predictions and management more difficult. Nevertheless, research is increasingly focused on understanding the causes and impacts of transient dynamics and finding ways to manage them. Examples include disease outbreaks Tao et al. (2021), shifts in species populations, and significant changes in ecosystems, such as coral reef degradation Norström et al. (2009). Early studies have also explored related concepts, such as cyclical plant succession and species’ behavior in lakes and savannas under varying conditions Van Langevelde et al. (2003); Scheffer et al. (1993). Recognizing the importance of transient dynamics is crucial because these temporary periods can lead to unexpected changes in ecosystems that might not be predicted by focusing solely on long-term behavior Hastings et al. (2018); Morozov et al. (2020). Understanding these dynamics is especially important when managing invasive species and in conservation and restoration efforts Hobbs and Harris (2001); Neubert and Caswell (2000); Hastings (2004); Shea and Chesson (2002); Hastings and Botsford (2006), where setting realistic recovery timelines and identifying critical points for intervention are essential. Additionally, transient dynamics reveal complex interactions between species, such as predator-prey relationships and competition, that may be less apparent once the ecosystem reaches a stable state Scheffer et al. (2005); Murdoch et al. (2003); Hastings (2001); Schreiber and Rudolf (2008); Holt (2008).
The importance and complexity of transient dynamics in physical systems have also recently gained attention. There have been reports of transient chaotic behavior in dynamical systems Dudkowski et al. (2022), even in cases where the attractor is hidden within the phase space Nag Chowdhury and Ghosh (2020). In fact, it may be possible that a dynamical system may become temporarily trapped at a local maximum or minimum within a potential landscape for a long time, giving the impression of stability. However, a minor disturbance—such as noise or an external force—can enable the system to overcome this barrier and transition to the lowest energy (or most stable) configuration Koch et al. (2024); Nag Chowdhury et al. (2020a); Bovier and Den Hollander (2016); Kelso (2012). From a different perspective, one recent study Parastesh et al. (2022) explores the role of higher-order interactions in neuronal dynamics by examining a simplicial complex of neurons, revealing that weak second-order interactions can facilitate synchronization at lower first-order coupling strengths, and that three-body interactions reduce overall synchronization costs compared to pairwise interactions. Similarly, Ref. Mehrabbeik et al. (2023a) investigates higher-order interactions in a memristive Rulkov model network, using master stability functions to analyze synchronization patterns, and demonstrates that incorporating higher-order interactions lowers the required coupling parameters for synchronization while also showing that larger network sizes enhance synchronization dynamics and facilitate cluster synchronization under specific coupling conditions. Many other intriguing studies on higher-order interactions Alvarez-Rodriguez et al. (2021); Pal et al. (2024); Ma et al. (2024); Anwar et al. (2024); Mirzaei et al. (2022); Mehrabbeik et al. (2023b); Chatterjee et al. (2023a); Vasilyeva et al. (2021); Alvarez-Rodriguez et al. (2022); Majhi et al. (2022); Kumar et al. (2021) exist; however, most of them primarily emphasis on long-term behaviors.
@@
On the other hand, traditionally, models have focused on either pairwise interactions Ray et al. (2022); Nag Chowdhury et al. (2024); Ray et al. (2020); Nag Chowdhury et al. (2023a); Chatterjee and Zehmakan (2023); Nag Chowdhury et al. (2019a) or higher-order interactions (HOI) Battiston et al. (2020); Gibbs et al. (2022); Zhang et al. (2023); Battiston et al. (2021) exclusively. However, real-world ecosystems are likely to exhibit a mix of these interaction types, with species interactions containing varying fractions of higher-order components. Such mixed-interaction frameworks are essential to better reflect the complexities of natural systems and provide more accurate insights into their dynamics. Recent research has focused on identifying HOIs from ecological data and making HOIs more common in ecological systems Terry et al. (2020); Li et al. (2021); Raj et al. (2024); Kleinhesselink et al. (2022); Barbosa et al. (2023). In classical niche theory Chase and Leibold (2009); Soberón and Nakamura (2009), species coexistence within the same niche is often challenged by competitive exclusion, leading weaker species to be out-competed by stronger ones. Conversely, neutral theory Brokaw and Busing (2000) posits that species are ecologically equivalent, with biodiversity emerging from a stochastic balance between speciation and extinction. While both theories have significantly advanced our understanding, they also have notable limitations. For instance, niche theory struggles with the paradox of the plankton, where the number of coexisting species exceeds the number of available limiting resources Huisman and Weissing (1999). Similarly, neutral theory’s “neutral drift” is incompatible with the observed stability in forest diversity Clark and McLachlan (2003).
@@
Recent developments in ecological modeling have introduced game theory-based approaches Nag Chowdhury et al. (2023b); Park et al. (2017); Nagatani et al. (2019); Nag Chowdhury et al. (2021a); Roy et al. (2022, 2023); Nagatani (2019); Ahmadi et al. (2023); Nag Chowdhury et al. (2021b); Yang and Park (2023), such as intransitive cyclic competition models like the rock-paper-scissors (RPS) game, which maintain biodiversity through cyclical dominance. These models demonstrate that even in hierarchical systems, intransitive competition can stabilize species coexistence, a phenomenon known as the “stabilizing effect of intransitivities.” Empirical evidence supports the prevalence of intransitive competition in various ecological communities, such as bacterial strains Kirkup and Riley (2004); Laird and Schamp (2006); Hibbing et al. (2010); Szolnoki et al. (2014); Park et al. (2017); Bhattacharyya et al. (2020); Islam et al. (2022); Chatterjee et al. (2023b) and phytoplankton Soliveres et al. (2015) and parasite-grass-forb Cameron et al. (2009). Although research on systems that consider both pairwise and higher-order interactions is still emerging, it is gaining attention Gibbs et al. (2022); Malizia et al. (2024); Kundu and Ghosh (2022). One recent study Van Giel et al. (2024) examined how the speed at which higher-order interactions emerge affects the stability and evolution of ecological networks, offering new insights into this underexplored area. Another study Gibbs et al. (2024) found that while higher-order interactions can lead to equilibrium in species abundance, they do not always ensure stable coexistence. However, when weak or cooperative pairwise interactions align with higher-order interactions, they can foster robust coexistence in diverse ecosystems. Using the generalized Lotka-Volterra model, Ref. Singh and Baruah (2021) derive a rule showing that while negative higher-order interactions can stabilize species coexistence by strengthening intraspecific competition, positive higher-order interactions can do so across a wider range of conditions by alleviating pairwise competition, and their results extend to multispecies communities, emphasizing the role of negative intraspecific HOIs in maintaining diversity. Our focus differs as we seek to better understand transient dynamics and explore how they respond to the simultaneous presence of both pairwise and higher-order interactions.
@@
Figure 1: The impact of pairwise interactions ($\gamma_{1}$) and higher-order interactions ($\gamma_{2}$) on species coexistence. Dynamics of model systems with N = 3 for (a) Pairwise $(\gamma_{1}=1.0)$ (b) Non-pairwise $(\gamma_{2}=1.0)$ (c) Mixed system $(\gamma_{1}\neq 0$, $\gamma_{2}\neq 0$, $\gamma_{1}+\gamma_{2}=1.0)$. The initial condition $(0.45,0.3,0.25)$ is chosen randomly within the $(0,1)\times(0,1)\times(0,1)$ maintaining the constraint $x_{1}(0)+x_{2}(0)+x_{3}(0)=1.0$. As shown in subfigure (a), the pairwise interaction system does not converge to a steady state. However, introducing higher-order interactions significantly alters this behavior. Interestingly, the system with both pairwise and higher-order interactions converges more slowly compared to the system with only higher-order interactions, although both ultimately reach the same equilibrium point $\big{(}\frac{1}{3},\frac{1}{3},\frac{1}{3}\big{)}$. $\gamma_{1}=\gamma_{2}=0.5$ is set for subfigure (c).
@@ II.1 Description
where $B_{ij}^{1}=A_{ij}-A_{ji}$, $B_{ijk}^{2}=2A_{ij}A_{ik}-A_{ji}A_{jk}-A_{ki}A_{kj},\text{ and }B_{ijkl}^{3}=3A_{ij}A_{ik}A_{il}-A_{ji}A_{jk}A_{jl}-A_{ki}A_{kj}A_{kl}-A_{li}A_{lj}A_{lk}\;\forall i,j=1,2,3,\cdots,N$.
Here, $\gamma_{i}\in[0,1]$ represents the contribution factor, which indicates the percentage of $B^{i}$ that contributes to the total interaction. So,
@@
$A$ matrix encodes information about interactions between two species. As there are $N$ species, $A$ is an $N\times N$ matrix. $A_{jk}$ term account for the winning probability of $j$-th species over the $k$-th species. Hence,
@@ II.2 Solution and Global Stability Analysis
Figure 2: The influence of varying proportions of pairwise interactions ($\gamma_{1}$) and higher-order interactions ($\gamma_{2}$) on species dynamics. Dynamics of mixed systems with $N=3$ for different values of $\gamma_{1}$ and $\gamma_{2}$: (a) $\gamma_{1}=0.7,\gamma_{2}=0.3$, (b) $\gamma_{1}=0.5,\gamma_{2}=0.5$, (c) $\gamma_{1}=0.3,\gamma_{2}=0.7$. Starting from the initial condition of $(0.45,0.3,0.25)$, as in Fig. 1, the system consistently converges to the coexistence equilibrium point, regardless of the relative contributions of pairwise and higher-order interactions. However, as the proportion of pairwise interactions decreases (i.e., as $\gamma_{1}$ diminishes), the transient time significantly shortens. This demonstrates that higher-order interactions play a crucial role in accelerating convergence towards equilibrium.
@@ II.2.1 Pairwise System
Now, we take the following Lyapunov function Chatterjee et al. (2022),
@@
Since $\sum_{j=1}^{N}A_{ij}x_{j}^{*}=\frac{1}{2}$ Grilli et al. (2017), we have
@@ II.2.3 Mixed System
$-\gamma_{2}\bigg{[}2\sum_{i=1}^{N}x_{i}^{*}\sum_{j=1}^{N}\sum_{k=1}^{N}\big{(}A_{ij}A_{ik}+A_{ij}A_{jk}\big{)}x_{j}x_{k}+1\bigg{]}$
Now, using the same simplification as mentioned in the above two analyses, we finally get for $\gamma_{2}\neq 0$,
$\frac{dV}{dt}=-\gamma_{1}\big{[}0\big{]}-\gamma_{2}\bigg{[}2\sum_{i=1}^{N}x_{i}^{*}\bigg{(}\sum_{j=1}^{N}A_{ij}\eta_{j}\bigg{)}^{2}\bigg{]}<0.$
@@ II.3 Transient dynamics
Figure 3: Reduction of transient time with increasing higher-order interactions ($\gamma_{2}$).The boxplot illustrates how transient times vary with different values of $\gamma_{2}$ across various initial conditions. For the considered interaction matrix $A$, we use ecologically relevant initial conditions and plot the transient times across varying levels of higher-order interactions. For a fixed value of $\gamma_{2}$, we select 20 distinct initial conditions and determine the transient times required to reach the coexistence equilibrium. We then calculate the arithmetic mean of these transient times, represented by the red dots, and the median, indicated by the yellow line. As $\gamma_{2}$ increases from $0.05$ (indicating a minimal presence of higher-order interactions) to $1.0$ (representing a system solely dominated by higher-order interactions), the transient times steadily decrease. While the higher values of $\gamma_{2}$ may appear to have minimal variation in transient times due to the large range of the y-axis, the inset provides a clearer view, showing that transient times with different initial conditions are actually well-distributed across this range.
Figure 4: Variation of (a) mean transient time and (b) the real part of the leading eigenvalue with respect to $\gamma_{2}$. (a) Mean transient time plotted against the proportion of higher-order interactions $\gamma_{2}$, with a rectangular hyperbola fit (red dashed line) illustrating how transient time decreases as $\gamma_{2}$ increases. (b) The most negative real part of the eigenvalues of the Jacobian matrix associated with the linearized system at the coexistence equilibrium plotted against $\gamma_{2}$ (blue dots), with a linear fit (red dashed line). The negative real part of the leading eigenvalue in (b) provides a direct measure of convergence speed. The correlation between the most negative real part of the leading eigenvalue and the mean transient time illustrates that greater negativity in the leading eigenvalue is associated with faster convergence to the same coexistence equilibrium point. This highlights the role of higher-order interactions in accelerating system stabilization and reducing transient oscillations.
@@ II.3.2 Role of Initial Abundances in Systems with a Fixed Interaction Matrix
In Figure 3, transient times are summarized using a boxplot, with the median indicated by the yellow line and the mean by the red dot. Connecting lines between adjacent mean values provide a clearer view of the trend. As $\gamma_{2}$ increases, indicating a higher proportion of higher-order interactions, mean transient times decrease. This suggests that ecosystems with more complex interactions can stabilize more rapidly, which is crucial for predicting how real-world ecosystems maintain biodiversity. An inset in the figure highlights the reduction in transient time, showing how quickly systems approach stability with higher-order interactions. To create this figure, we selected $20$ different random initial conditions, each constrained by $x_{1}(0)+x_{2}(0)+x_{3}(0)=1.0$. We measured the transient time for each condition—defined as the period until density oscillations fall below $10^{-6}$ and then calculated the average of these transient times to obtain the mean value.
When $\gamma_{2}=0$, the system’s transient time is effectively infinite, leading to a rapid decline near zero and a more gradual decrease as $\gamma_{2}$ increases. To quantify this relationship, we fitted the mean transient time for $\gamma_{2}$ values ranging from $0.05$ to $1.0$ with a rectangular hyperbola fit in Fig. 4(a), illustrating how the product of mean transient time and $\gamma_{2}$ remains constant, say $c$, as $\gamma_{2}$ increases. Our statistical analysis suggests that the value of this constant is $c=343.1173$ with a Squared Error Loss of $2705.41$. The Squared Error Loss can be further reduced by increasing the number of initial data points, thereby providing a larger dataset for fitting. This fit provides a statistical measure of how swiftly the system’s transient time diminishes as the proportion of higher-order interactions increases. Note that the fitted inverse relationship between $\gamma_{2}$ and the mean transient time indicates that as $\gamma_{2}$ approaches $0^{+}$, the mean transient time tends to infinity, which aligns with our observations. Moreover, the fit demonstrates that the transient duration decreases as $\gamma_{2}$ increases, but this reduction occurs in a non-linear fashion. This finding highlights the significant impact of higher-order interactions on accelerating the system’s convergence to equilibrium.